OpenStudy (anonymous):

Use the Remainder Theorem to find the remainder resulting from the division: (x^6-8)÷(x-√2)

OpenStudy (triciaal):

x^6 -8 = (x^2 -2)(x^4 + x^2*2 + 4 )= (x^2-2)(x^2 + 2)^2 - x^2*2 if x-rt2 is a factor when x = rt 2 the remainder will be 0 if not a factor when x= rt 2 the value is the remainder